Find the x-intercepts of the graph of the equation:
y = x2 - 1

Answer by stanbon(75887)

Find the x-intercepts of the graph of the equation:
y = x2 - 1
----
Let y = 0.
Solve: x^2-1 = 0
Factor:
(x-1)(x+1) = 0
x = 1 or x = -1
----
x-intercepts: (1,0) and (-1,0)
===============================
Cheers,
Stan H.